# EKF–SINDy: Empowering the extended Kalman filter with sparse identification of nonlinear dynamics

### v1

@@ 3.2.1 Time-delay embedding
In general, (fully) learning $\mathbf{f}$ directly in its original dimension from lower dimensional measurements $\mathbf{y}$ is not possible. To tackle this issue, time-delay embedding techniques art:Brunton17 can be adopted to lift the low-dimensional time series into a high-dimensional space where it is possible to recover hidden state components that might be fundamental to describe the dynamics bakarji2023discovering . In our context, we hope to recover features carrying on information about the unobserved oscillator.
Firstly, the time-delay embedding method is applied and the associated Hankel matrix is constructed by stacking time-shifted copies of the observed time series. Secondly, a Singular Value Decomposition (SVD) of this matrix is performed to extract time-delay coordinates approximating the Koopman operator in finite dimension art:Dylewsky22 , thereby providing a linearised description of the system dominant dynamics art:Champion19 . Indeed, the reference system formed by these time-delay coordinates is diffeomorphic to the original attractor of the system dynamics, under the conditions given by the Takens’ embedding theorem proc:Takens81 . Lastly, this new set of coordinates can be leveraged by SINDy both in the training and prediction phases, to better capture the dynamics and to improve prediction performance bakarji2023discovering .
Putting the method into practice, we report the outcome of the time-delay embedding including in the Hankel matrix the dependence on the stiffness $k_{2}$ of the unobserved oscillator, thus setting $\phi=k_{2}$. These embedded coordinates will be exploited to construct the SINDy model. For the $\iota$-th sample $\phi_{\iota}$, with $\iota=1,\ldots,I$, Eqs. (3.2) are numerically integrated in time for $T\leq T$. It is worth noting that, unlike what has been done here, the procedure does not strictly require to collect snapshots over the whole time axis. By collecting $z_{1}(\phi_{\iota},t_{j})$, for $j=1,\ldots,T$, we assemble the matrix:
@@
Thus, we construct the Hankel matrix $\mathbf{A}$ by stacking $\mathbf{A}=[\mathbf{A}_{1},\ldots,\mathbf{A}_{I}]\in\mathbb{R}^{w\times(T-w)I}$, and next we take its SVD:
$\mathbf{A}=\mathbf{U}\mathbf{S}\mathbf{V}^{\top},$ (22)
where: $\mathbf{U}\in\mathbb{R}^{w\times w}$ and $\mathbf{V}\in\mathbb{R}^{(T-w)I\times(T-w)}I$ are two orthogonal matrices whose columns are termed left and right singular vectors, respectively; ${\mathbf{S}}\in\mathbb{R}^{w\times(T-w)}I$ is a pseudo-diagonal matrix collecting the associated singular values.
For this case, we have considered $I=16$, $w=200$ and $\zeta=1$, where $w$ and $\zeta$ indicate the number of delay embeddings and the embedding period (ı.e, the number of lags in each successive embedding), respectively. The $I$ values of the parameter $k_{2}$ have been sampled in the range $[1.0,2.0]$ using a stratified sampling book:Saltelli07 .
Performing the SVD on the Hankel matrix indicates that only $4$ dominant modes are responsible for the $99.97\%$ of the variance of $\mathbf{A}$, as illustrated by Fig. 5. This outcome aligns with the problem original dimension.In Fig. 5, we have also highlighted that the left singular values (reported in the boxes close to the corresponding singular values) resemble Legendre polynomials, as expected by choosing a small embedding period ($\zeta=1$) vautard1989singular ; broomhead1986extracting .
Coming back to the time delay parameters $w$ and $d$, we can argue that our choices guaranteed to unfold the attractor in the embedding space ma2006selection by satisfying the Taken’s embedding theorem condition on the number of delay embeddings $w>2n$. Indeed, this condition guarantees the existence of the diffeomorphism to the original system. A comparison of time-delayed coordinates with respect to the original (unobserved) ones is reported in Fig. 5. The similarity between these coordinates highlights that the time-delay coordinates adopted effectively allow for unfolding the attractor.
According to the singular value decay, we have truncated the SVD matrices to account just for the $n=4$ dominant modes. The selected singular values and the left singular vectors are collected in the matrices $\tilde{\mathbf{S}}\in\mathbb{R}^{n\times n}$ and $\tilde{\mathbf{U}}\in\mathbb{R}^{w\times n}$, respectively.
Figure 5: Nonlinear resonator. Dominant singular values and left singular vectors (left) and corresponding time-delay coordinates (right), obtained by performing the SVD on the Hankel matrix constructed by integrating the considered nonlinear dynamic system for $I=16$ values of $k_{2}$ sampled from $[1.0,2.0]$.
@@ 3.2.2 EKF-SINDy leveraging embedded coordinates
The state vector is then augmented with the corresponding parameter as $\boldsymbol{\varkappa}=\left[\mathbf{x}^{\top},\phi\right]^{\top}$. State data are collected in the matrix $\mathbf{X}_{\varkappa}\in~{}\mathbb{R}^{TI\times(n+1)}$ that is necessary for training the SINDy model. According to Algorithm 1, we use SINDy to build the function $\mathbf{f}\approx\boldsymbol{\Xi}_{f}\boldsymbol{\Theta}_{f}(\boldsymbol{\varkappa})$, describing the dynamics of the delayed coordinates. Polynomial terms up to the third order are included in $\boldsymbol{\Theta}_{f}$, due to the cubic nonlinearities present in the system (3.2). The sparsity promoting term has been set to $L=10^{-3}$. Interestingly, diminishing this regularisation term to $L=5\cdot 10^{-4}$ including more terms in the description of the dynamics, slightly improves the reconstruction capacity of SINDy but greatly deteriorates the identification capacity of the procedure. This suggests that the extra term included in the description of the system dynamics obtained for $L=5\cdot 10^{-4}$ does not have a physical meaning, but simply helps in fitting the collected time histories art:Champion19 .
On the other hand, the function $\mathbf{h}$ extracting observations $z_{1}(t_{j})$ from the state variables $\mathbf{x}(t_{j})$ is known. Specifically, $\mathbf{h}$ is linear and it can be determined by inverting Eq. (23) as in the following
@@ 3.2.3 Estimation of the stiffness $k_{2}$ of the hidden oscillator
Figure 6: Nonlinear dynamic system, estimation of the resonator stiffness $\bar{k}=1.2$. In the top left corner, the nonlinear resonator model is depicted, with the resonator stiffness to be calibrated highlighted in red. In the top right corner, the estimated parameter $\hat{k}_{2}$ is plotted in red against the target $\bar{k}_{2}$ in black. The red shaded area represents the $95\%$ confidence interval of the estimates, determined using the posterior covariance. In the other graphs, the evolution of the (hidden) system respoonse tracked by the filter is plotted in red against the real system dynamics in black.
@@
We have applied the EKF-SINDy procedure to estimate the stiffness $k_{2}$ of the hidden oscillator. In the offline phase, we have performed time-delay embedding and approximate the dynamic model $\mathbf{f}$ by training SINDy on the time-delayed coordinates. The parameters defining the operated time delay-embedded have been reported in Sec. 3.2.1. As previously specified, $k_{2}$ has been sampled from $[1.0,2.0]$.
@@
We test our method for two instances $\bar{k}_{2}\in\{1.2,2.3\}$, whose results are reported in Fig. 6 and Fig. 8, respectively. The tuning parameters of the filter are reported in the Appendix.
@@
For the test case $\bar{k}_{2}=1.2$, starting from an initial parameter guess outside the training range of SINDy and underestimating $\bar{k}_{2}$ by $25\%$, the EKF-SINDy method progressively increases accuracy and decreases uncertainty until $t=~{}50$, when it converges to the correct value. Similarly, it provides an accurate state estimate with uncertainty bounds that include the (partially observed) state of the system. We recall that data assimilation is limited to the observation of the displacement of the first oscillator, while the other quantities required to describe the system ($z_{2}$, $\dot{z}_{1}$ and $\dot{z}_{2}$) are unobserved. The target $\bar{\mathbf{x}}=[\bar{x}_{1},\bar{x}_{2},\bar{x}_{3},\bar{x}_{4}]^{\top}$ is reconstructed from the real dynamics of the system for sake of comparison with the filter estimates $\hat{\mathbf{x}}_{n}$. After convergence at $t=~{}50$, the predicted mean value of the state closely matches the evolution of the state and demonstrates robustness over the entire time horizon. The reconstructed $\hat{z}_{1}$ is plotted against the acquired measurements $\bar{z}_{1}$ in Fig. 7. demonstrating the noise tolerance of the procedure, with the filter estimate $\hat{z}_{1}$ closely aligning with the noise-free version of the signal $\hat{\bar{z}}_{1}$.
@@
Figure 7: Nonlinear dynamic system, estimation of the resonator stiffness $\bar{k}=1.2$. Comparison between the observed variable $\bar{z}_{1}$ in black and its filter reconstruction $\hat{z}_{1}$ in red. The noise free version $\hat{\bar{z}}_{1}$ of the acquired measurements is depicted with a dotted line in pink.
@@
Starting the identification process from an initial guess for $k_{2}$ outside the limits for which data have been collected in the offline phase, as shown in Algorithm 1, assesses the robustness of procedure. Operating outside such a training range is hardly achievable by other ML techniques like NNs chp:SpringerChp22 ; art:CAS22 , which constitutes a major advantage of the proposed procedure. This advantage stems from the capacity of the retained SINDy library terms to describe the dominant dynamical behaviour of the system. Thus, these terms approximate the physical process that underlies the observed dynamics art:Brunton16 .
@@
To further evaluate the capability of the proposed approach to function beyond the training range of SINDy, a second test case with $\bar{k}_{2}=2.3$ has been conducted. In this case, the initial guess for the stiffness of the second oscillator overestimated the correct value by $20\%$. Unlike the scenario with $\bar{f}_{2}=1.2$, the entire identification procedure was expected to operate outside the training range $[1.0,2.0]$ of SINDy. However, EKF-SINDy successfully performed the estimation of the system properties as illustrated in Fig. 8. In contrast, the estimation of the hidden state components $x_{3}$ and $x_{4}$ shows some degradation.
@@
Figure 8: Nonlinear dynamic system, estimation of the resonator stiffness $\bar{k}=2.3$. See caption of Fig. 6 for further details.
@@ 3.2.4 Estimation of the parameters $\alpha$ and $\beta$ related to linear and quadratic coupling terms
To demonstrate the procedure capability of identifying more than one system parameter, we have simultaneously addressed the estimation of the linear and quadratic coupling terms $\alpha$ and $\beta$, while setting $k_{2}=1.95f_{1}$. In future work, we will investigating the internal resonances potentially arising for this frequency ratio as done in art:Giorgio21 .
We have sampled $I=20$ values of $\alpha$ and $\beta$ to construct the Hankel matrix and to train SINDy on the time-delayed coordinates. Specifically, we have applied the Latin hypercube sampling across the domain $[0.005,0.5]^{2}$ with logarithmic scaling book:Saltelli07 . The logarithmic scaling has been used to put attention to cases featuring, possibly simultaneously, low values of $\alpha$ and $\beta$. Indeed, considering only high values of $\alpha$ and $\beta$ will possibly make the system response dynamics ruled by few dominant dynamic effects, precluding the investigation of potential interaction phenomena. As in previous cases, we have fixed the number $w$ of delay embeddings to $200$, and the embedding period $\zeta$ at $1$. The tuning of the filter is reported in the Appendix.
@@
Figure 9: Nonlinear dynamic system, estimation of the parameters $\alpha$ and $\beta$ of the linear and quadratic coupling terms. In the top left corner, the nonlinear resonator model is depicted with highlighted the coupling parameters to be calibrated. In the top right corner, the estimated parameter $\hat{\alpha}$ is plotted in red against the target $\bar{\alpha}$ in black. Below, the estimated $\hat{\beta}$ is plotted in blue against the target $\bar{\beta}$ in black. The red and blue shaded area represents $95\%$ confidence interval of the estimates, determined using the posterior covariance. In the bottom right corner, a comparison between the observed variable $\bar{z}_{1}$ in black and its filter reconstruction $\hat{z}_{1}$ in red is reported. The noise free version $\hat{\bar{z}}_{1}$ of the acquired measurements is depicted with a dotted line in pink. A zoom in of this plot is reported on the left.
Figure 10: Nonlinear dynamic system, estimation of the parameters $\alpha$ and $\beta$ of the linear and quadratic coupling terms. The evolution of (hidden) system dynamics tracked by the filter is plotted in red against the real dynamics in black. The red shaded area represents $95\%$ confidence interval of the estimates determined using posterior covariance. For the sake of presentation, the outcomes are truncated at $t=80$.

### v2

@@ 3.2.1 Time-delay embedding
In general, (fully) learning $\mathbf{f}$ directly in its original dimension from lower dimensional measurements is not possible using SINDy, as it requires knowledge of the full underlying state space variables [27]. To tackle this issue, we adopt the time-delay embedding techniques [10] to lift the low-dimensional time series $y(t)=z_{1}(t)$ into a high-dimensional space (see Sec. 2.2.2). With time-delay embedding is possible to recover hidden features carrying on information about the unobserved oscillator, which might be fundamental to describe the system’s dynamics [11].
Putting the method into practice, we report the outcome of the time-delay embedding including in the Hankel matrix the dependence on the stiffness $k_{2}$ of the unobserved oscillator, thus setting $\phi=k_{2}$. These embedded coordinates will be exploited to construct the SINDy model. For the $\iota$-th sample $\phi_{\iota}$, with $\iota=1,\ldots,I$, Eqs. (24) are numerically integrated in time. By collecting $z_{1}(\phi_{\iota},t_{j})$, for $j=1,\ldots,T$, we assemble the matrix:
@@
Thus, we construct the Hankel matrix $\mathbf{A}$ by stacking $\mathbf{A}=[\mathbf{A}_{1},\ldots,\mathbf{A}_{I}]\in\mathbb{R}^{w\times(T-w)I}$, and next we take its SVD $\mathbf{A}=\mathbf{U}\mathbf{S}\mathbf{V}^{\top},$ with $\mathbf{U}\in\mathbb{R}^{w\times w}$ and $\mathbf{V}\in\mathbb{R}^{(T-w)I\times(T-w)I}$, such that $\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}$ and $\mathbf{V}^{\top}\mathbf{V}=\mathbf{I}$, while ${\mathbf{S}}\in\mathbb{R}^{w\times(T-w)}I$ is a pseudo-diagonal matrix collecting the associated singular values $s_{\lambda}$, with $\lambda=1,\ldots,w$.
For this case, we have considered $I=16$, $w=200$ and $\zeta=1$, where $w$ and $\zeta$ indicate the number of delay embeddings and the the number of lags in each successive embedding, respectively. The $I$ values of the parameter $k_{2}$ have been sampled in the range $[1.0,4.0]$ using a stratified sampling [62]. In practice, the range $[1.0,4.0]$ has been partitioned into $I$ subintervals, and random sampling has been performed from each subinterval. The number of parameter values required to detect parametric dependence is related to the complexity of this dependency and the problem at hand. In this specific case, being a didactic example, we chose a number that is sufficiently large for this purpose. In applications where the computation of system trajectories can be extremely costly in terms of computational resources, a more parsimonious choice of the number of instances $I$ can be considered.
Performing the SVD on the Hankel matrix indicates that only $4$ dominant modes are responsible for the $99.97\%$ of the variance of $\mathbf{A}$, as illustrated by Fig. 6. This outcome aligns with the problem original dimension. In Fig. 6, we have also highlighted that the left singular values (reported in the boxes close to the corresponding singular values) resemble Legendre polynomials, as expected by choosing a small embedding period ($\zeta=1$) [63, 64].
Coming back to the time delay parameters $w$ and $\zeta$, we can argue that our choices guaranteed to unfold the attractor in the embedding space [43] by satisfying the Taken’s embedding theorem condition on the number of delay embeddings $w>2n$. Indeed, this condition guarantees the existence of the diffeomorphism to the original system. A comparison of time-delayed coordinates with respect to the original (unobserved) ones is reported in Fig. 6. The similarity between these coordinates highlights that the time-delay coordinates adopted effectively allow for unfolding the attractor. According to the singular value decay, we have truncated the SVD matrices to account just for the $n=4$ dominant modes. The selected singular values and the left singular vectors are collected in the matrices $\tilde{\mathbf{S}}\in\mathbb{R}^{n\times n}$ and $\tilde{\mathbf{U}}\in\mathbb{R}^{w\times n}$, respectively.
Figure 6: Nonlinear resonator. Dominant singular values $s_{\lambda}$ and left singular vectors (left) and corresponding time-delay coordinates (right), obtained by performing the SVD on the Hankel matrix constructed by integrating the considered nonlinear dynamical system for $I=16$ values of $k_{2}$ sampled from $[1.0,4.0]$.
@@ 3.2.2 EKF-SINDy leveraging embedded coordinates
The state vector is then augmented with the corresponding parameter as $\boldsymbol{\varkappa}=\left[\mathbf{x}^{\top},\phi\right]^{\top}$. State data are collected in the matrix $\mathbf{X}_{\varkappa}\in~{}\mathbb{R}^{TI\times(n+1)}$ that is necessary for training the SINDy model. According to Algorithm 1, we use SINDy to build the function $\mathbf{f}\approx\boldsymbol{\Xi}_{f}\boldsymbol{\Theta}_{f}(\boldsymbol{\varkappa})$, describing the dynamics of the delayed coordinates. Polynomial terms up to the third order are included in $\boldsymbol{\Theta}_{f}$, due to the cubic nonlinearities present in the system (24). Sparsity is promoted using STLSQ with $L=10^{-3}$. Interestingly, diminishing this threshold term to $L=5\cdot 10^{-4}$ including more terms in the description of the dynamics, slightly improves the reconstruction capacity of SINDy but greatly deteriorates the identification capacity of the procedure. This suggests that the extra term included in the description of the system dynamics obtained for $L=5\cdot 10^{-4}$ does not have a physical meaning, but simply helps in fitting the collected time histories [27]. On the other hand, the function $\mathbf{h}$ extracting observations $z_{1}(t_{j})$ from the state variables $\mathbf{x}(t_{j})$ is known. Specifically, $\mathbf{h}$ is linear and it can be determined by inverting Eq. (26) as in the following
@@ 3.2.3 Estimation of the stiffness $k_{2}$ of the hidden oscillator
Figure 7: Nonlinear dynamical system, estimation of the resonator stiffness $\bar{k}_{2}=1.44$. In the top graph, the estimated parameter $\hat{k}_{2}$ is plotted in red against the target $\bar{k}_{2}$ in black. The red shaded area represents the $95\%$ confidence interval of the estimates, determined using the posterior covariance. In the other graphs, the evolution of the (hidden) system response tracked by the filter is plotted in red against the real system dynamics in black.
@@
We have applied the EKF-SINDy procedure to estimate the stiffness $k_{2}$ of the hidden oscillator. In the offline phase, we have performed time-delay embedding and approximate the dynamic model $\mathbf{f}$ by training SINDy on the time-delayed coordinates. The parameters defining the operated time delay-embedded have been reported in Sec. 3.2.1. As previously specified, $k_{2}$ has been sampled from $[1.0,4.0]$.
@@
We test our method for two instances $\bar{k}_{2}\in\{1.44,5.29\}$, whose results are reported in Fig. 7 and Fig. 9, respectively. The tuning parameters of the filter are reported in Appendix B.
@@
For the test case $\bar{k}_{2}=1.44$, starting from an initial parameter guess outside the training range of SINDy and underestimating $\bar{k}_{2}$ by $35\%$, the EKF-SINDy method progressively increases accuracy and decreases uncertainty until $t=50$, when it converges to the correct value. Similarly, it provides an accurate state estimate with uncertainty bounds that include the (partially observed) state of the system. We recall that data assimilation is limited to the observation of the displacement of the first oscillator, while the other quantities required to describe the system ($z_{2}$, $\dot{z}_{1}$ and $\dot{z}_{2}$) are unobserved. The target $\bar{\mathbf{x}}=[\bar{x}_{1},\bar{x}_{2},\bar{x}_{3},\bar{x}_{4}]^{\top}$ is reconstructed from the real dynamics of the system for sake of comparison with the filter estimates $\hat{\mathbf{x}}_{n}$. After convergence at $t=~{}50$, the predicted mean value of the state closely matches the evolution of the state and demonstrates robustness over the entire time horizon. The reconstructed $\hat{z}_{1}$ is plotted against the acquired measurements $\bar{z}_{1}$ in Fig. 8. demonstrating the noise tolerance of the procedure, with the filter estimate $\hat{z}_{1}$ closely aligning with the noise-free version of the signal $\hat{\bar{z}}_{1}$.
@@
Figure 8: Nonlinear dynamical system, estimation of the resonator stiffness $\bar{k}=1.44$. Comparison between the observed variable $\bar{z}_{1}$ in black and its filter reconstruction $\hat{z}_{1}$ in red. The noise free version $\hat{\bar{z}}_{1}$ of the acquired measurements is depicted with a dotted line in pink.
@@
Starting the identification process from an initial guess for $k_{2}$ outside the limits for which data have been collected in the offline phase, as shown in Algorithm 1, assesses the robustness of procedure. Operating outside such a training range is hardly achievable by other ML techniques like NNs [65, 66], which constitutes a major advantage of the proposed procedure. This advantage stems from the capacity of the retained SINDy library terms to describe the dominant dynamical behaviour of the system. Thus, these terms approximate the physical process that underlies the observed dynamics [5].
@@
To further evaluate the capability of the proposed approach to function beyond the training range of SINDy, a second test case with $\bar{k}_{2}=7.62$ has been conducted. In this case, the initial guess for the stiffness of the second oscillator overestimated the correct value by $20\%$. Unlike the scenario with $\bar{k}_{2}=1.2$, the entire identification procedure was expected to operate outside the training range $[1.0,2.0]$ of SINDy. However, EKF-SINDy successfully performed the estimation of the system properties as illustrated in Fig. 9. In contrast, the estimation of the hidden state components $x_{3}$ and $x_{4}$ shows some degradation.
@@
Figure 9: Nonlinear dynamical system, estimation of the resonator stiffness $\bar{k}=5.29$. See caption of Fig. 7 for further details.
@@ 3.2.4 Estimation of the parameters $\alpha$ and $\beta$ related to linear and quadratic coupling terms
To demonstrate the procedure capability of identifying more than one system parameter, we have simultaneously addressed the estimation of the linear and quadratic coupling terms $\alpha$ and $\beta$, while setting $k_{2}=1.95^{2}k_{1}$. In future work, we will investigate the internal resonances potentially arising for this ratio as done in [61].
We have sampled $I=20$ values of $\alpha$ and $\beta$ to construct the Hankel matrix and to train SINDy on the time-delayed coordinates. Specifically, we have applied the Latin hypercube sampling across the domain $[0.005,0.5]^{2}$ defined for $\alpha$ and $\beta$ with logarithmic scaling [62]. The logarithmic scaling has been used to put attention to cases featuring, possibly simultaneously, low values of $\alpha$ and $\beta$. Indeed, considering only high values of $\alpha$ and $\beta$ will possibly make the system response dynamics ruled by few dominant dynamic effects, precluding the investigation of potential interaction phenomena. As in previous cases, we have fixed the number $w$ of delay embeddings to $200$, and the embedding period $\zeta$ at $1$. The tuning of the filter is reported in Appendix B.
@@
Figure 10: Nonlinear dynamical system, estimation of the parameters $\alpha$ and $\beta$ of the linear and quadratic coupling terms. In the top left corner, the estimated parameter $\hat{\alpha}$ is plotted in red against the target $\bar{\alpha}$ in black. On the right, the estimated $\hat{\beta}$ is plotted in blue against the target $\bar{\beta}$ in black. The red and blue shaded area represents $95\%$ confidence interval of the estimates, determined using the posterior covariance. In the bottom right corner, a comparison between the observed variable $\bar{z}_{1}$ in black and its filter reconstruction $\hat{z}_{1}$ in red is reported. The noise free version $\hat{\bar{z}}_{1}$ of the acquired measurements is depicted with a dotted line in pink. A zoom in of this plot is reported on the left.
Figure 11: Nonlinear dynamical system, estimation of the parameters $\alpha$ and $\beta$ of the linear and quadratic coupling terms. The evolution of (hidden) system dynamics tracked by the filter is plotted in red against the real dynamics in black. The red shaded area represents $95\%$ confidence interval of the estimates determined using posterior covariance. For the sake of presentation, the outcomes are truncated at $t=80$.